# The Green Peace Dividend: the Effects of Militarization on Emissions and the Green Transition

### v1

@@ 4 The Green Peace Dividend
First of all, it is most straightforward to quantify the effect of military buildups on GHG emissions. This is what the empirical analysis and most of the model analysis was focused on. The empirical analysis indicates that a one percentage point rise in the military spending share leads to a rise in total emissions of between 1 and 1.5%, depending on the method and sample used. The model indicates, under the baseline calibration for the US in 2017, that a permanent one percentage point rise in the military spending share raises GHG emissions by between 0.36% and 1.81%, depending on the composition of military spending.
@@
Focusing solely on these direct effects of military buildups on GHG emissions, it is possible to translate the damages due to induced emissions into dollar terms using the social cost of carbon (SCC). For example, if one uses the social cost of carbon (SCC) provided by the EPA (2023), equal to \$190 per ton of CO2 at a near-term discount rate of 2%, it is possible to consider a scenario in which the US permanently doubled its military spending share in 2017 and to calculate the climate damages resulting from this policy change using the SCC. Total GHG emissions in the US in 2017 were equal to 6.09 billion tons of CO2e. The model estimate of a 1.18-6% rise in emissions in response to such a shock (with a military spending share of 3.3% in 2017) leads to global climate damages equivalent to approximately 13.6-69.4 billion dollars per year - or approximately 0.07-0.35% of 2017 US GDP. [^30] This calculation ignores the transition path to the new steady-state, as well as the fact that emissions change from their 2017 level for reasons independent of defence policy. More recent - and larger - estimates, such as the SCC calculated by Bilal and Känzig (2024), equal to \$1056 per ton of CO2, would yield an interval of 75.9-385.9 billion dollars per year - or approximately 0.39-2% of 2017 US GDP. Moreover, if, along with the US, other countries - such as NATO members, or Russia and China - increased their military spending share, costs would be even higher.
@@
Instead of looking at damages, one can reframe the consequences in the following way: rising GHG emissions due to military buildups lead to the quicker exhaustion of our carbon budget. [^31] The amount of GHG emissions we can afford before breaching a given threshold of warming, such as 1.5 degrees Celsius. Hence, militarization would reduce our chances of avoiding any given global warming limit.
These damage calculations omit several effects which cannot be properly quantified. First of all, both the empirical and the model results show that emission intensity increases significantly in response to a military buildup. The reason for this, as the model analysis showed, is that energy-intensive and fossil fuel industries experience an expansion. As argued previously, this can lead to a rise in their political influence and lobbying power, which could make passing carbon policies and the green transition much more complicated. It could also reduce green lobbying efforts, with the same outcome.
@@
Second, the model indicates that investments in energy-intensive industries crowd out investments in other sectors after a military spending shock. In particular, investment in the utilities sector slumps, which is worrying because this can translate into a slump in renewables investments. In addition, if R&D investments in this and other sectors are crowded out, then green innovation could also be scaled down. In turn, fossil fuel and energy-intensive industries could scale up "dirty" or "grey" R&D expenditures, which could contribute to a widening gap relative to clean technical progress - as Acemoglu et al. (2012) argued, such a gap could be difficult to overcome, and could be costly to close.
@@
To sum up, military buildups impact the carbon budget and cause climate damages directly by increasing GHG emissions. In addition, they might significantly complicate the green transition by crowding out green investment and R&D funds and by bolstering fossil-fuel and energy-intensive industries that could obstruct the passing and implementation of climate policies.
5 Policy implications
@@ 6 Conclusion
In this paper, I show that military buildups increase greenhouse gas emissions and can significantly complicate the green transition. The empirical analysis suggests that a one percentage point rise in the military spending share increases total emissions by 1-1.5%, and emission intensity by 1%.
The dynamic production network model, calibrated using US data from 2017, indicates that military buildups increase both total emissions, as well as emission intensity. The magnitude of these effects, however, seems to be lower than the one indicated by the empirical analysis. Permanently doubling the military spending share raises total emissions by between 0.36% and 1.81%, while it raises emission intensity by between 0.22% and 1.5%.
Hence, military buildups can directly reduce our carbon budget. However, the model analysis also indicates that militarization can impede the green transition by crowding out investment in renewables and, possibly, green R&D. Similarly, it leads to a considerable expansion of fossil fuel and energy-intensive industries, which, by consequently increasing their political influence and reducing green lobbying, might complicate passing and implementing carbon policies.

### v3

@@ 4 The Green Peace Dividend and policy implications
First of all, it is most straightforward to quantify the effect of military buildups on GHG emissions. This is what the empirical analysis and most of the model analysis was focused on. The empirical analysis indicates that a one percentage point rise in the military spending share leads to a rise in total emissions of between 0.9 and 2%, depending on the method and sample used. The model indicates, under the baseline calibration for the US in 2017, that a permanent one percentage point rise in the military spending share raises GHG emissions by between 0.36% and 1.81%, depending on the composition of military spending.
@@
I use three main methods to assess the impacts of the measured effects on GHG: by using the social cost of carbon (SCC), by using the carbon budget, and by looking at the global path of surface temperatures.
@@
First, focusing solely on these direct effects of military buildups on GHG emissions, it is possible to translate the damages due to induced emissions into dollar terms using the social cost of carbon (SCC). For example, if one uses the social cost of carbon (SCC) provided by the EPA (2023), equal to \$190 per ton of CO2 at a near-term discount rate of 2%, it is possible to consider a scenario in which the US permanently doubled its military spending share in 2017 and to calculate the climate damages resulting from this policy change using the SCC. Total GHG emissions in the US in 2017 were equal to 6.09 Gt of CO2e. The model estimate of a 1.18-6% rise in emissions in response to such a shock (with a military spending share of 3.3% in 2017) leads to global climate damages equivalent to approximately 13.6-69.4 billion dollars per year - or approximately 0.07-0.35% of 2017 US GDP. [^32] This calculation ignores the transition path to the new steady-state, as well as the fact that emissions change from their 2017 level for reasons independent of defence policy. More recent - and larger - estimates, such as the SCC calculated by Bilal and Känzig (2024), equal to \$1056 per ton of CO2, would yield an interval of 75.9-385.9 billion dollars per year - or approximately 0.39-2% of 2017 US GDP. Note that these simple calculations ignore the convexity of climate damages.
@@
Second, one can reframe the consequences in the following way: rising GHG emissions due to military buildups lead to the quicker exhaustion of our carbon budget. [^33] The amount of GHG emissions we can afford before breaching a given threshold of warming, such as 2 degrees Celsius. Hence, militarization would reduce our chances of avoiding any given global warming limit. In particular, according to the Global Carbon Budget 2023 (Friedlingstein et al. (2023)), the remaining carbon budget for a 50% likelihood to limit global warming to 2 °C is equal to 1150 Gt of CO2 from the beginning to 2024. Under 2023 emission levels (approximately 40.9 Gt of CO2), this would be exhausted in 28 years. Now, assuming that all countries permanently raise their military expenditure share by 3.3 percentage points (as we assumed in the SCC calculations for the US), and that the effects are similar to the US (emissions rise by 1.18-6%), then the carbon budget could be exhausted 0.22-1.5 years earlier.
Third, one can check how a global military buildup of 3.3 percentage points would change the path of temperatures, under different shared socioeconomic pathways (SSP) (Riahi et al. (2017)). Here, I use the central estimate of the effect of a military expenditure shock on emissions, which implies that a permanent shock of 3.3 p.p. raises GHG emissions by 3%. Using live.magicc.org, an online software relying on MAGICC7 (Meinsshausen et al. (2020)) (which is an integrated assessment model), I calculate that the difference would be minimal under the SSP1-1.9 scenario (which assumes that GHG emissions are very low, and are reduced to zero by 2050), with temperatures being higher by approximately 0.01 °C in 2100 (see figure 17). [^34] In the simulations, I set CO2 emissions from fossil fuel use and industrial activity to be higher by 3%, after 2015. However, under the SSP3 scenario (where GHG emissions double by 2100), temperatures would be higher by 0.05 °C by 2100, indicating that the effect of a military buildup becomes worse under worse scenarios. Under the SSP5 scenario, temperatures would be higher by more than 0.06 °C. [^35] For an explanation of the different shared socioeconomic pathways, see: https://www.carbonbrief.org/explainer-how-shared-socioeconomic-pathways-explore-future-climate-change/ This indicates that under progressively worse scenarios, the effect of a military buildup on temperatures also becomes stronger, as a 3% rise in a higher baseline amount of emissions leads to a larger change in the level of emissions, leading, in turn, to a larger change in the level of the temperature. Note that temperatures in 2100 rise with the shown scenarios (so they are highest under SSP5, and lowest under SSP1 - 1.9), and temperatures peak before 2100 only under the SSP1 scenarios.
Figure 17: Effects on temperature under different SSP scenarios
Note: the figure shows by how much higher surface temperatures would be in 2100 given that a permanent global military buildup of 3.3 p.p. happens, in °C, under different shared socioeconomic pathways.
These damage calculations omit several effects which cannot be properly quantified. First of all, the empirical results show that green patenting falls significantly following a military spending shock. This could, in itself, retard the transition to net zero emissions. However, if fossil fuel and energy-intensive industries do not reduce or even scale up "dirty" or "grey" R&D expenditures, this could contribute to a widening gap relative to clean technical progress - as Acemoglu et al. (2012) argued, such a gap could be difficult to overcome, and could be costly to close.
@@
Second, the model indicates that investments in energy-intensive industries crowd out investments in other sectors after a military spending shock. In particular, investment in the utilities sector slumps under a temporary military spending shock, which is worrying because this can translate into a slump in renewables investments, as the majority of investments in this sector are directed towards renewables.
Third, both the empirical and the model results show that emission intensity increases significantly in response to a military buildup. The reason for this, as the model analysis showed, is that energy-intensive and fossil fuel industries experience an expansion. As argued previously, this can lead to a rise in their political influence and lobbying power, which could make passing carbon policies and the green transition much more complicated.
Also, note that if there indeed is a causal effect of climate change on the risk of conflict, then a positive feedback loop generated by the impact of military buildups on climate change increases the damages due to both of these phenomena.
@@
To sum up, military buildups impact the carbon budget and cause climate damages directly by increasing GHG emissions. In addition, they might significantly complicate the green transition by crowding out green innovation and investment and by bolstering fossil-fuel and energy-intensive industries that could obstruct the passing and implementation of climate policies.
Policy
@@ 5 Conclusion
In this paper, I show that military buildups increase greenhouse gas emissions and can significantly complicate the green transition. The empirical analysis suggests that a one percentage point rise in the military spending share increases total emissions by 0.9-2%, and emission intensity by 1%. It also suggests that such a shock can reduce green patenting by 10-25%.
The dynamic production network model, calibrated using US data from 2017, indicates that military buildups increase both total emissions, as well as emission intensity. The magnitude of these effects, however, seems to be lower than the one indicated by the empirical analysis. Permanently doubling the military spending share raises total emissions by between 0.36% and 1.81%, while it raises emission intensity by between 0.22% and 1.5%. Hence, military buildups can directly reduce our carbon budget. However, the model analysis also indicates that militarization can impede the green transition by crowding out investment in renewables, in the case of a temporary military spending shock.